OpenStudy (anonymous):

Solve the following system of equations using substitution. What is the value of y? 2x+3y=105 x+2y=65 I somehow got -1(4/6)x=-5. Can someone provide a clear explanation on how to do this?

OpenStudy (texaschic101):

x + 2y = 65 ==> x = -2y + 65 now sub -2y + 65 in for x in the other equation 2x + 3y = 105 2(-2y + 65) + 3y = 105 -- distribute -4y + 130 + 3y = 105 -- subtract 130 -4y + 3y = 105 - 130 -- combine like terms -y = - 25 y = 25 x + 2y = 65 x + 2(25) = 65 x + 50 = 65 x = 65 - 50 x = 15 check.. x + 2y = 65 2x + 3y = 105 15 + 2(25) = 65 2(15) + 3(25) = 105 15 + 50 = 65 30 + 75 = 105 65 = 65 (correct) 105 = 105 (correct) I know you just needed to find y, but I found them both in order to check them. y = 25 and x = 15

OpenStudy (texaschic101):

do you understand, or do you have any questions ?

OpenStudy (anonymous):

I get it know, thanks a bunch :)

OpenStudy (texaschic101):

anytime :)